Circular No. 7927

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dr IN UGC 12214 D. Hutchings and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 17.0) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on June 24.5 and 25.5 UT. SN 2002dr is located at α = 22^{h}51^{m}01^{s}.00, δ = +31^{o}22′23′′.0 (equinox 2000.0), which is 0′′.5 west and 7′′.1 south of the nucleus of UGC 12214. A KAIT image taken on 2001 Dec. 13.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2002do, 2002dp, 2002dq T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra (range 370–750 nm) were obtained by R. Cho and M. Hicken on June 19.43 (SN 2002do; cf. IAUC 7923), 20.42 (SN 2002dq; cf. IAUC 7924), and 20.47 UT (SN 2002dp; cf. IAUC 7924) with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph). SNe 2002do and 2002dp are both type-Ia supernovae near maximum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocities of 4761 and 3489 km/s for the host galaxies of SN 2002do and 2002dp, respectively, the supernova expansion velocities are ∼ 10300 and ∼ 11300 km/s for Si II (rest 635.5 nm). The spectral-feature ages are 1 ± 2 days before maximum light for both SNe 2002do and 2002dp. SN 2002dq is a type-II supernova whose spectrum consists of a blue continuum with hydrogen P-Cyg lines; adopting the NED recession velocity of 2519 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 11200 km/s.

COMET C/2002 C1 (IKEYA-ZHANG) Ephemeris extension to IAUC 7901 (elements on MPC 45654):

2002 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 25 15^{h}22.^{m}46 + 3^{o}20.^{′}4 1.125 1.944 130.^{o}4 23.^{o}5  8.6
July 5 15 22.19 − 0 43.9 1.348 2.091 123.7 23.9  9.2
15 15 24.69 − 3 58.2 1.585 2.234 116.6 24.0  9.8
25 15 29.25 − 6 37.3 1.835 2.373 109.3 23.8 10.3
Aug. 4 15 35.34 − 8 50.5 2.093 2.510 102.0 23.3 10.8
14 15 42.63 −10 44.3 2.358 2.643  94.8 22.4 11.2
24 15 50.85 −12 22.9 2.626 2.775  87.6 21.4 11.6
Sept. 3 15 59.82 −13 49.0 2.895 2.903  80.5 20.0 12.0
2002 June 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green